OpenStudy (laronjaxon):

Please help, In the following equation, identify the x-intercepts in the graph y = 3x2 + 19x – 40

OpenStudy (mertsj):

(3x-5)(x+8)=0 3x-5=0 or x+8=0 x=5/3, or x = -8